# One Token Can Help! Learning Scalable and Pluggable Virtual Tokens for Retrieval-Augmented Large Language Models

### v3

@@ Abstract
Retrieval-augmented generation (RAG) is a promising way to improve large language models (LLMs) for generating more factual, accurate, and up-to-date content. Existing methods either optimize prompts to guide LLMs in leveraging retrieved information or directly fine-tune LLMs to adapt to RAG scenarios. Although fine-tuning can yield better performance, it often compromises the LLMs’ general generation capabilities by modifying their parameters. This limitation poses challenges in practical applications, especially when LLMs are already deployed, as parameter adjustments may affect their original functionality. To address this, we propose a novel method that involves learning scalable and pluggable virtual tokens for RAG. By maintaining the LLMs’ original parameters and fine-tuning only the embeddings of these pluggable tokens, our approach not only enhances LLMs’ performance but also preserves their general generation capabilities. Furthermore, we design several training strategies to improve the scalability, flexibility, and generalizability of our method. Comprehensive experiments across nine question-answering tasks demonstrate the superiority of our approach. The virtual tokens and training code are publicly available at https://github.com/DaoD/SPRING/.
One Token Can Help! Learning Scalable and Pluggable Virtual Tokens for Retrieval-Augmented Large Language Models
Yutao Zhu, Zhaoheng Huang, Zhicheng Douthanks: Correpsonding author.,  and Ji-Rong Wen Gaoling School of Artificial Intelligence, Renmin University of China yutaozhu94@gmail.com, dou@ruc.edu.cn
1 Introduction
@@
Large language models (LLMs) have achieved remarkable performance across various natural language processing tasks Brown et al. (2020); OpenAI (2023); Touvron et al. (2023). Despite their extensive parameters enabling them to learn rich knowledge during pre-training, LLMs may still generate hallucinated, outdated, or inaccurate content, especially in scenarios requiring long-tail knowledge Ji et al. (2023); Zhang et al. (2023b). To address this problem, retrieval-augmented generation (RAG) has emerged as a pivotal strategy. By explicitly decoupling knowledge retrieval from the backbone LLMs, such architectures have achieved more accurate and reliable content generation and shown particularly enhanced performance on knowledge-intensive tasks such as open-domain question answering Lewis et al. (2020b); Petroni et al. (2021); Izacard et al. (2023); Tan et al. (2024); Jin et al. (2024).
@@
Existing efforts in RAG development can be roughly categorized into two groups (as illustrated in Figure 1). The first group leverages the in-context learning capabilities of LLMs by incorporating retrieved information into the input along with appropriate prompts Shi et al. (2023); Ram et al. (2023). This allows for straightforward application to any off-the-shelf LLM without tuning its parameters. However, its effectiveness largely depends on the human experience in crafting effective prompts and the LLM’s ability to interpret these prompts. The second group focuses on training LLMs to enhance their performance in RAG scenarios. This training might involve either end-to-end pre-training Guu et al. (2020); Borgeaud et al. (2022) or fine-tuning Lin et al. (2023); Wang et al. (2023) for specific tasks. These approaches can often lead to better performance, but they require significant computational resources. Recently, parameter-efficient fine-tuning techniques, such as LoRA Hu et al. (2022), have been widely studied, significantly reducing training costs. These methods can optimize the LLMs’ parameters for RAG, but unfortunately compromise the model’s general generation abilities in scenarios without retrieval. This limitation prevents their application to LLMs already operational in real-world settings.
@@
In experiments, SPRING is trained with Mistral-7b, LLaMA-2-7b, and LLaMA-2-13b on nine commonly used QA datasets and evaluated in both in-domain and out-of-domain tasks. The experimental results demonstrate that SPRING not only effectively improves the RAG performance of LLMs but also successfully preserves their general generation capabilities. Overall, the SPRING method exhibits four main characteristics:
$\bullet$ Lightweight yet effective. Instead of updating the full parameters of the LLMs, we opt to freeze the pre-trained models and only learn the embeddings for the added virtual tokens. For example, adding 50 tokens to the Mistral-7b model introduces only 0.2M parameters in total. Despite these minimal parameters, SPRING improves the average EM and F1 scores by more than 33% and 12% across nine QA datasets, respectively.
@@
2 Related Work
Retrieval-Augmented Generation (RAG)
@@ 2 Related Work
Compared to standard text generation, RAG incorporates a retrieval module that accesses external knowledge to enhance generation quality Lewis et al. (2020b); Guu et al. (2020); Zhu et al. (2023). The mainstream RAG follows a “retrieve-then-read” paradigm, where the retrieval module provides external knowledge as additional context, which is then read by generation models to produce the final output Izacard et al. (2023); Shi et al. (2023); Ram et al. (2023); Borgeaud et al. (2022); Lin et al. (2023); Zhu et al. (2024). To optimize the use of external knowledge, some methods focus on crafting effective prompts that guide the utilization of retrieved information Shi et al. (2023); Ram et al. (2023). These prompt-based methods are applicable to any LLM without tuning its parameters. However, they depend heavily on skillful prompt writing and the LLMs’ ability to understand instructions. In contrast, other studies attempts to directly train the model to better use the retrieved knowledge. For example, REALM Guu et al. (2020) and RETRO Borgeaud et al. (2022) incorporate retrieval in end-to-end retrieval-augmented pre-training. RA-DIT Lin et al. (2023) employs fine-tuning to enhance LLMs’ retrieval understanding. These tuning-based methods often yield better performance than prompt-based methods by optimizing model parameters for RAG. However, they may compromise the LLMs’ general capabilities, particularly in non-retrieval scenarios. Different from existing methods, we design a new lightweight tuning method for RAG. It is a plug-and-play module that enhances RAG performance using trainable virtual tokens, which can be removed in non-RAG scenarios to preserve the LLMs’ general generation abilities.
Parameter-Efficient Fine-Tuning (PEFT)
@@
The paradigms of “pre-training then fine-tuning” have demonstrated efficacy across various natural language Devlin et al. (2019); Raffel et al. (2020); Lewis et al. (2020a); Radford et al. (2019) and vision tasks He et al. (2020); Dosovitskiy et al. (2021); Chen et al. (2020). The common fine-tuning process involves tuning all parameters of a model, which is computational intensive, especially for LLMs. To address this, PEFT Mangrulkar et al. (2022) approaches have been developed. These approaches freeze most of the pre-trained models’ parameters, yet still manage to achieve comparable performance on downstream tasks. PEFT has been widely studied Wan et al. (2023), and typical methods including adapter-based tuning Houlsby et al. (2019); Lin et al. (2020); Rebuffi et al. (2017); Chen et al. (2023), low-rank adaptation (LoRA) Hu et al. (2022); Dettmers et al. (2023), and prompt tuning Li and Liang (2021); Lester et al. (2021); Liu et al. (2021b, a); Qin and Eisner (2021). Adapter-based tuning inserts lightweight modules into a model’s existing layers and have been extended to various domains Gao et al. (2024); Hu et al. (2023); Zhang et al. (2023a). LoRA Hu et al. (2022) introduces trainable low-rank matrices that adjust the model’s weight updates, achieving promising fine-tuning performance on LLMs Hu et al. (2023). Prompt tuning incorporates a series of trainable prompt tokens to LLMs. These tokens can be inserted either to the input layer only Lester et al. (2021); Liu et al. (2021b) or to all of the intermediate layers Li and Liang (2021); Liu et al. (2021a). In this paper, we proposes a novel prompt tuning method, SPRING, specifically designed for RAG scenarios. Our method introduces virtual tokens between retrieved results and the input, exploiting the auto-regressive generation paradigm to improve the model’s ability to utilize retrieved information. Additionally, it is designed to be scalable and pluggable, thus broadening its application scope while preserving the original generative capabilities of LLMs.
@@ 3.1 Problem Formulation
$p_{\text{LM}}=\prod_{i=1}^{n}p_{\theta}(x_{i}|x_{<i}),$ (1)
where $x_{<i}$ denotes the sequence of tokens preceding $x_{i}$ at each step, and $\theta$ represents the parameters of the model. For RAG, a retrieval corpus $\mathcal{D}$ and a retriever ${M}$ are introduced. Then, the generation process is conditioned on both $x_{x<i}$ and the retrieved results ${R}={M}_{\mathcal{D}}(x_{x<i})$ as:
$p_{\text{RAG}}$ $=\prod_{i=1}^{n}p_{\theta}(x_{i}|{R};x_{<i}),$ (2)
$p_{\text{RAG-QA}}$ $=\prod_{i=1}^{n}p_{\theta}(a_{i}|{R};Q).$ (3)
@@
3.2 SPRING: Scalable and Pluggable Virtual Tokens for RAG
Our SPRING method, shown in Figure 2, introduces trainable virtual tokens into the input to optimize LLMs for RAG scenarios. Specifically, following the notation in Equation (3), we add $n$ trainable tokens $T=[t_{1},t_{2},\cdots,t_{n}]$ between the retrieved results ${R}_{\mathcal{D}}(Q)$ and the input $Q$ [^1] In our implementation, we set $n=50$ for training. The generation process can then be described as:
$p_{\text{{SPRING}{}}}=\prod_{i=1}^{n}p_{\theta,\delta}(a_{i}|{R};[t_{1},t_{2},\cdots,t_{n}];Q),$ (4)
@@ 3.2 SPRING: Scalable and Pluggable Virtual Tokens for RAG
Importantly, we place the virtual tokens $T$ between the retrieved results $R$ and the question $Q$ for two main reasons: (1) In the auto-regressive generation paradigm, positioning the tokens after the retrieved results allows them to attend to this information, thereby aiding the model’s comprehension. (2) Recent studies have indicated that LLMs are particularly sensitive to the end of an input Liu et al. (2023). By consistently placing these virtual tokens before the question across all test samples, we aim to mitigate any potential adverse effects on the understanding of the question.
@@
In practical developments, LLMs are often constrained by their maximum input lengths, limiting the number of tokens available for retrieval augmentation (especially when the retrieved results are very long). Therefore, it is desired to design a mechanism so that any number of virtual tokens can be used in the inference to improve RAG performance. To achieve this, we propose an optimization strategy working like a “spring” (as shown in Figure 2). Specifically, for a given sample $\{R,Q,A\}$ with the total number of added tokens $n$, we randomly select a number $k(k\leq n)$ and utilize the first $k$ virtual tokens $t_{1:k}$ to construct the training example as $[R;t_{1},t_{2},\cdots,t_{k};Q]$. This method allows for the flexible optimization of any number of virtual tokens. Consequently, the number of virtual tokens incorporated during inference can be dynamically adjusted based on the requirements of the application. The effectiveness of this strategy and its comparison with other methods are further discussed in Section 4.4.2.
@@
We illustrate the instructions for using our SPRING in Figure 3. After training, the embeddings of the added tokens have been optimized for RAG, but these tokens do not correspond to any existing tokens in the vocabulary. To make them easy to use, we can add some special tokens (e.g., [r1], $\cdots$, [r50]) to the vocabulary and initialize their embeddings with the trained embeddings. Then, during inference, after obtaining the retrieved results $R$, we can add any number of these special tokens (e.g., [r1] $\cdots$ [rk]) after $R$ and input them with the question $Q$ to the LLMs for generation. We also provide an example code snippet in Appendix A for using our method in practice.
We refer to our method as SPRING due to its scalable and pluggable nature, making it particularly well-suited for enhancing existing LLMs that are already deployed. Additionally, it effectively bridges the gap between retrieved results and user input, significantly improving the LLMs’ capabilities in understanding the retrieved external knowledge.
Table 1: Evaluation results of different backbone models and methods on nine QA datasets. The retriever is e5-large model, and the number of retrieved passages is set as three. The number of virtual tokens used in SPRING is set as 50. PopQA is invisible during training. “Prefix” stands for prefix-tuning. The best results are in bold.
@@
Dataset Metric Concat Prompt Prefix LoRA SPRING Concat Prompt Prefix LoRA SPRING
Tuning Parameters 0 0 0.2M 4M 0.2M 0 0 0.2M 4M 0.2M
TQA EM 0.00 57.79 11.74 62.76 65.71 0.01 39.90 0.00 0.03 46.56
F1 65.60 80.33 59.97 85.44 85.26 63.96 69.72 28.30 34.91 74.48
NQ EM 0.00 28.99 13.04 47.95 42.35 0.00 13.36 0.00 0.00 18.80
F1 41.77 58.72 38.22 74.15 70.73 43.74 48.63 17.74 29.10 55.75
HQA EM 0.00 26.36 5.79 39.95 35.26 0.00 17.07 0.00 0.03 20.15
F1 44.91 56.15 42.59 68.93 65.44 47.54 49.50 17.45 27.57 54.79
SQuAD EM 0.00 23.92 7.19 35.71 33.67 0.00 8.61 0.00 0.00 12.71
F1 43.05 57.66 39.75 68.05 66.99 43.32 46.81 21.88 27.51 53.58
WebQ EM 0.00 17.53 4.44 43.65 31.84 0.00 14.79 0.00 0.00 24.95
F1 37.46 52.10 31.55 71.99 64.78 44.34 50.60 20.14 32.36 59.83
2Wiki EM 0.00 22.64 4.38 35.93 31.80 0.00 23.45 0.00 0.01 24.62
F1 47.77 55.58 41.82 63.85 62.03 52.83 53.55 21.14 37.09 56.83
CoQA EM 0.00 8.20 1.56 12.89 13.28 0.00 8.59 0.00 0.00 9.96
F1 27.98 36.72 20.02 41.19 42.41 32.97 36.58 13.15 18.99 39.96
MS MARCO EM 0.00 5.73 0.60 8.13 6.57 0.00 2.56 0.00 0.01 2.09
F1 56.44 53.81 50.56 54.81 53.48 49.50 47.75 47.44 52.44 51.41
PopQA EM 0.00 39.79 10.02 47.15 48.71 0.00 16.05 0.00 0.00 20.25
F1 56.49 68.26 44.54 73.12 73.90 53.61 54.85 20.39 25.09 58.32
Average EM 0.00 25.66 6.53 37.13 34.35 0.00 16.04 14.31 0.01 20.01
F1 46.83 57.70 41.00 66.84 65.00 47.98 50.89 23.07 31.67 56.11
4 Experiment
4.1 Datasets and Retrievers
@@
We conduct experiments on nine commonly used question-answering datasets, including TriviaQA (TQA) Joshi et al. (2017), Natural Questions (NQ) Kwiatkowski et al. (2019), HotpotQA (HQA) Yang et al. (2018), SQuAD 1.1 Rajpurkar et al. (2016), Web Questions (WebQ) Berant et al. (2013), 2WikiMultiHopQA (2Wiki) Ho et al. (2020), CoQA Reddy et al. (2019), MS MARCO Nguyen et al. (2016), and PopQA Mallen et al. (2023). These datasets are publicly available at HuggingFace or their official websites. Since PopQA only has a test set, we use it as a held-out dataset to evaluate the generalizability of the methods. Except for PopQA, we mix the training set of all other datasets for training, and test the methods on their corresponding test set. If the test set is unavailable, we use the development set for evaluation. It is worth noting that, though some datasets have provided golden reference passages for the answer, we do not use them in our experiment but use the passages retrieved from the following retrieval sets in both training and inference stages. Exact match (EM) and F1 score are employed as evaluation metrics.
@@
For the retrieval sets, we follow previous studies Yoran et al. (2023) and use the combination of Wikipedia and MS MARCO datasets as the retrieval corpus. Wikipedia contains high-quality human knowledge, which is helpful for many knowledge-intensive tasks. MS MARCO contains a large amount of web pages, which can provide information necessary for curating some natural language questions. We use the datasets that have already been preprossed into passages and released on HuggingFace. [^2] Wikipedia passages: https://huggingface.co/datasets/Tevatron/wikipedia-nq-corpus. MS MARCO passages: https://huggingface.co/datasets/Tevatron/msmarco-passage-corpus. The Wikipedia set has 21M passages, while the MS MARCO set has 8M passages. More details are provided in Appendix B.
@@
We use E5-large Wang et al. (2022) as the main retriever in our experiments. The impact of other retrievers, i.e., BM25 Robertson and Zaragoza (2009), BGE-base Xiao et al. (2023), and E5-base, is studied in our further analysis (Section 4.4.3). Among these retrievers, BM25 is a non-neural sparse retrieval algorithm, while others are neural-based dense retrievers. In general, dense retrievers perform better than BM25 on several benchmarks Wang et al. (2022); Izacard et al. (2022).
4.2 Baseline Methods
@@
We consider both the base and instruction-tuned versions of Mistral-7b, LLaMA-2-7b, and LLaMA-2-13b as the backbone models, and compare our SPRING with the following baselines.

### v4

@@ Abstract
Retrieval-augmented generation (RAG) is a promising way to improve large language models (LLMs) for generating more factual, accurate, and up-to-date content. Existing methods either optimize prompts to guide LLMs in leveraging retrieved information or directly fine-tune LLMs to adapt to RAG scenarios. Although fine-tuning can yield better performance, it often compromises the LLMs’ general generation capabilities by modifying their parameters. This limitation poses challenges in practical applications, especially when LLMs are already deployed, as parameter adjustments may affect their original functionality. To address this, we propose a novel method that involves learning scalable and pluggable virtual tokens for RAG. By maintaining the LLMs’ original parameters and fine-tuning only the embeddings of these pluggable tokens, our approach not only enhances LLMs’ performance but also preserves their general generation capabilities. Furthermore, we design several training strategies to improve the scalability, flexibility, and generalizability of our method. Comprehensive experiments across 12 question-answering tasks demonstrate the superiority of our approach.
Introduction
@@
Large language models (LLMs) have achieved remarkable performance across various natural language processing tasks (Brown et al. 2020; OpenAI 2023; Touvron et al. 2023). Despite their extensive parameters enabling them to learn rich knowledge during pre-training, LLMs may still generate hallucinated, outdated, or inaccurate content, especially in scenarios requiring long-tail knowledge (Ji et al. 2023; Zhang et al. 2023b). To address this problem, retrieval-augmented generation (RAG) has emerged as a pivotal strategy. By explicitly decoupling knowledge retrieval from the backbone LLMs, such architectures have achieved more accurate and reliable content generation and shown particularly enhanced performance on knowledge-intensive tasks such as open-domain question answering (Petroni et al. 2021; Tan et al. 2024; Jin et al. 2024b).
@@
Existing efforts in RAG development can be roughly categorized into two groups (as illustrated in Figure 1). The first group leverages the in-context learning capabilities of LLMs by incorporating retrieved information into the input along with appropriate prompts (Shi et al. 2023; Ram et al. 2023). This allows for straightforward application to any off-the-shelf LLM without tuning its parameters. However, its effectiveness largely depends on the human experience in crafting effective prompts and the LLM’s ability to interpret these prompts. The second group focuses on training LLMs to enhance their performance in RAG scenarios. This training might involve either end-to-end pre-training (Guu et al. 2020; Borgeaud et al. 2022) or fine-tuning (Lin et al. 2023; Wang et al. 2023b) for specific tasks. These approaches can often lead to better performance, but they require significant computational resources. Recently, parameter-efficient fine-tuning techniques, such as LoRA (Hu et al. 2022), have been widely studied, significantly reducing training costs. These methods can optimize the LLMs’ parameters for RAG, but unfortunately compromise the model’s general abilities in non-RAG scenarios, such as commonsense reasoning and in-context learning. All these limitations prevent their application to LLMs already operational in real-world settings.
@@
In experiments, SPRING is trained with the base and instruction fine-tuned versions of Mistral-7b, LLaMA-2-7b, and LLaMA-2-13b models and evaluated on 12 commonly used QA datasets, covering both in-domain and out-of-domain scenarios. The experimental results demonstrate that SPRING not only effectively improves the RAG performance of LLMs but also successfully preserves their general generation capabilities. Overall, the SPRING method exhibits four main characteristics:
$\bullet$ Lightweight yet effective. Instead of updating the full parameters of the LLMs, we opt to freeze the pre-trained models and only learn the embeddings for the added virtual tokens. For example, adding 50 tokens to the Mistral-7b model introduces only 0.2M parameters in total. Despite these minimal parameters, SPRING improves the average EM and F1 scores by more than 43% and 17% across 12 QA datasets, respectively.
@@
Related Work
Retrieval-Augmented Generation
@@
Compared to standard text generation, retrieval-augmented generation (RAG) incorporates a retrieval module that accesses external knowledge to enhance generation quality (Lewis et al. 2020; Guu et al. 2020; Zhu et al. 2023; Jin et al. 2024a). The mainstream RAG follows a “retrieve-then-read” paradigm, where the retrieval module provides external knowledge as additional context, which is then read by generation models to produce the final output (Shi et al. 2023; Ram et al. 2023; Borgeaud et al. 2022; Lin et al. 2023; Zhu et al. 2024). To optimize the use of external knowledge, some methods focus on crafting effective prompts that guide the utilization of retrieved information (Shi et al. 2023; Ram et al. 2023). These prompt-based methods are applicable to any LLM without tuning its parameters. However, they depend heavily on skillful prompt writing and the LLMs’ ability to understand instructions. In contrast, other studies attempts to directly train the model to better use the retrieved knowledge. For example, REALM (Guu et al. 2020) and RETRO (Borgeaud et al. 2022) incorporate retrieval in end-to-end retrieval-augmented pre-training. RA-DIT (Lin et al. 2023) employs fine-tuning to enhance LLMs’ retrieval understanding. These tuning-based methods often yield better performance than prompt-based methods by optimizing model parameters for RAG. However, they may compromise the LLMs’ general capabilities, particularly in non-retrieval scenarios. Different from existing methods, we design a new lightweight tuning method for RAG. It is a plug-and-play module that enhances RAG performance using trainable virtual tokens, which can be removed in non-RAG scenarios to preserve the LLMs’ general generation abilities.
Parameter-Efficient Fine-Tuning
@@
The paradigms of “pre-training then fine-tuning” have demonstrated efficacy across various natural language (Devlin et al. 2019; Raffel et al. 2020; Radford et al. 2019) and vision tasks (He et al. 2020; Dosovitskiy et al. 2021; Chen et al. 2020). The common fine-tuning process involves tuning all parameters of a model, which is computational intensive, especially for LLMs. To address this, parameter-efficient fine-tuning (PEFT) (Mangrulkar et al. 2022) approaches have been developed. These approaches freeze most of the pre-trained models’ parameters, yet still manage to achieve comparable performance on downstream tasks. PEFT has been widely studied (Wan et al. 2023), and typical methods including adapter-based tuning (Houlsby et al. 2019; Lin, Madotto, and Fung 2020; Rebuffi, Bilen, and Vedaldi 2017; Chen et al. 2023), low-rank adaptation (LoRA) (Hu et al. 2022; Dettmers et al. 2023), and prompt tuning (Li and Liang 2021; Lester, Al-Rfou, and Constant 2021; Liu et al. 2021b, a; Qin and Eisner 2021). Adapter-based tuning inserts lightweight modules into a model’s existing layers and have been extended to various domains (Gao et al. 2024; Hu et al. 2023; Zhang et al. 2023a). LoRA (Hu et al. 2022) introduces trainable low-rank matrices that adjust the model’s weight updates, achieving promising fine-tuning performance on LLMs (Hu et al. 2023). Prompt tuning incorporates a series of trainable prompt tokens to LLMs. These tokens can be inserted either to the input layer only (Lester, Al-Rfou, and Constant 2021; Liu et al. 2021b) or to all of the intermediate layers (Li and Liang 2021; Liu et al. 2021a). In this paper, we proposes a novel prompt tuning method, SPRING, specifically designed for RAG scenarios. Our method introduces virtual tokens between retrieved results and the input, exploiting the auto-regressive generation paradigm to improve the model’s ability to utilize retrieved information. Additionally, it is designed to be scalable and pluggable, thus broadening its application scope while preserving the original generative capabilities of LLMs.
@@
$p_{\text{LM}}=\prod_{i=1}^{m}p_{\theta}(x_{i}|x_{<i}),$ (1)
where $x_{<i}$ denotes the sequence of tokens preceding $x_{i}$ at each step, and $\theta$ represents the parameters of the model. For RAG, a retrieval corpus $\mathcal{D}$ and a retriever ${M}$ are introduced. Then, the generation process is conditioned on both $x_{<i}$ and the retrieved results ${R}={M}_{\mathcal{D}}(x_{<i})$ as:
$p_{\text{RAG}}$ $=\prod_{i=1}^{m}p_{\theta}(x_{i}|{R};x_{<i}),$ (2)
$p_{\text{RAG-QA}}$ $=\prod_{i=1}^{m}p_{\theta}(a_{i}|{R};Q;a_{<i}).$ (3)
@@
Scalable and Pluggable Virtual Tokens for RAG
Our SPRING method, shown in the left side of Figure 2, introduces trainable virtual tokens into the input to optimize LLMs for RAG scenarios. Specifically, following the notation in Equation (3), we add $n$ trainable tokens $T=[t_{1},t_{2},\cdots,t_{n}]$ between the retrieved results ${R}$ and the input $Q$. The generation process can then be described as:
$p_{\text{{SPRING}{}}}=\prod_{i=1}^{m}p_{\theta,\delta}(a_{i}|{R};[t_{1},t_{2},\cdots,t_{n}];Q;a_{<i}),$
@@
Importantly, we place the virtual tokens $T$ between the retrieved results $R$ and the question $Q$ for two main reasons: (1) In the auto-regressive generation paradigm, positioning the tokens after the retrieved results allows them to attend to this information, thereby aiding the model’s comprehension. (2) Recent studies have indicated that LLMs are particularly sensitive to the end of an input (Liu et al. 2023). By consistently placing these virtual tokens before the question across all test samples, we aim to mitigate any potential adverse effects on the understanding of the question.
@@
In practical developments, LLMs are often constrained by their maximum input lengths, limiting the number of tokens available for retrieval augmentation (especially when the retrieved results are very long). Therefore, it is desired to design a mechanism so that any number of virtual tokens can be used in the inference to improve RAG performance. To achieve this, we propose an optimization strategy working like a “spring” (as shown in Figure 2). Specifically, for a given sample $\{R,Q,A\}$ with the total number of added tokens $n$, we randomly select a number $k(k\leq n)$ and utilize the first $k$ virtual tokens $t_{1:k}$ to construct the training example as $[R;t_{1},t_{2},\cdots,t_{k};Q]$. This method allows for the flexible optimization of any number of virtual tokens. Consequently, the number of virtual tokens incorporated during inference can be arbitrarily selected based on the requirements of the application. The effectiveness of this strategy and its comparison with other methods are further discussed in our experiments.
@@
We illustrate the instructions for using our SPRING in the right side of Figure 2. After training, the embeddings of the added tokens have been optimized for RAG, but these tokens do not correspond to any existing tokens in the vocabulary. To make them easy to use, we can add some special tokens (e.g., [r1], $\cdots$, [r50]) to the vocabulary and initialize their embeddings with the trained embeddings. Then, during inference, after obtaining the retrieved results $R$, we can add any number of these special tokens (e.g., [r1] $\cdots$ [rk]) after $R$ and input them with the question $Q$ to the LLMs for generation. We also provide an example code snippet in Appendix for using our method in practice.
We refer to our method as SPRING due to its scalable and pluggable nature, making it particularly well-suited for enhancing existing LLMs that have already been deployed. Additionally, it effectively bridges the gap between retrieved results and user input, significantly improving the LLMs’ capabilities in understanding the retrieved external knowledge.
@@
Dataset Metric Concat Prompt Prefix LoRA SPRING SPRING+ Concat Prompt Prefix LoRA SPRING SPRING+
Tuning Parameters 0 0 0.2M 4M 0.2M 4.2M 0 0 0.2M 4M 0.2M 4.2M
Trivia QA EM 0.00 57.79 11.74 62.76 65.71 63.89 0.01 39.90 0.00 0.03 46.56 43.37
F1 65.60 80.33 59.97 85.44 85.26 85.94 63.96 69.72 28.30 34.91 74.48 74.89
NQ EM 0.00 28.99 13.04 47.95 42.35 49.78 0.00 13.36 0.00 0.00 18.80 25.54
F1 41.77 58.72 38.22 74.15 70.73 75.22 43.74 48.63 17.74 29.10 55.75 60.82
HQA EM 0.00 26.36 5.79 39.95 35.26 41.14 0.00 17.07 0.00 0.03 20.15 23.38
F1 44.91 56.15 42.59 68.93 65.44 69.95 47.54 49.50 17.45 27.57 54.79 57.99
SQuAD EM 0.00 23.92 7.19 35.71 33.67 35.98 0.00 8.61 0.00 0.00 12.71 13.90
F1 43.05 57.66 39.75 68.05 66.99 68.41 43.32 46.81 21.88 27.51 53.58 54.93
WebQ EM 0.00 17.53 4.44 43.65 31.84 48.10 0.00 14.79 0.00 0.00 24.95 28.81
F1 37.46 52.10 31.55 71.99 64.78 73.88 44.34 50.60 20.14 32.36 59.83 62.95
2Wiki EM 0.00 22.64 4.38 35.93 31.80 37.12 0.00 23.45 0.00 0.01 24.62 28.60
F1 47.77 55.58 41.82 63.85 62.03 64.60 52.83 53.55 21.14 37.09 56.83 59.12
CoQA EM 0.00 8.20 1.56 12.89 13.28 13.87 0.00 8.59 0.00 0.00 9.96 12.50
F1 27.98 36.72 20.02 41.19 42.41 42.04 32.97 36.58 13.15 18.99 39.96 41.36
MS MARCO EM 0.00 5.73 0.60 8.13 6.57 8.27 0.00 2.56 0.00 0.01 2.09 3.24
F1 56.44 53.81 50.56 54.81 53.48 55.90 49.50 47.75 47.44 52.44 51.41 49.84
PopQA* EM 0.00 39.79 10.02 47.15 48.71 46.98 0.00 16.05 0.00 0.00 20.25 18.70
F1 56.49 68.26 44.54 73.12 73.90 73.29 53.61 54.85 20.39 25.09 58.32 58.05
Fermi* EM 0.00 0.06 0.00 0.12 0.18 0.24 0.00 0.06 0.06 0.00 0.06 0.19
F1 21.66 18.16 12.65 29.63 31.15 30.67 25.51 17.84 20.33 25.42 29.28 30.37
Musique* EM 4.01 4.05 0.04 10.86 8.80 12.58 0.00 1.93 0.66 0.17 3.64 4.43
F1 38.79 38.64 20.91 52.65 48.74 53.65 42.30 36.18 35.22 45.31 44.93 47.70
Bamboogle* EM 12.80 12.80 0.00 24.22 22.66 28.13 0.00 4.69 3.20 0.00 12.00 12.80
F1 45.81 48.13 13.67 59.05 56.95 62.00 42.61 42.24 36.22 45.14 47.69 46.89
Average EM 1.40 19.80 4.90 30.78 28.40 32.17 0.00 12.59 0.33 0.02 16.32 17.96
F1 43.98 51.30 34.69 61.91 60.15 62.96 45.19 46.19 24.95 33.41 52.24 53.74
Table 1: Evaluation results of different methods on twelve QA datasets. The retriever is E5-large model, and the number of retrieved passages is set as three. The number of virtual tokens used in SPRING is set as 50. PopQA, Fermi, Musique, and Bamboogle are invisible during training. “Prefix” stands for prefix-tuning, and “SPRING+” is trained based on the LoRA’s checkpoint. The best results are in bold.
Experiment
Datasets and Retrievers
@@
We conduct experiments on twelve commonly used question-answering datasets, including TriviaQA (TQA) (Joshi et al. 2017), Natural Questions (NQ) (Kwiatkowski et al. 2019), HotpotQA (HQA) (Yang et al. 2018), SQuAD 1.1 (Rajpurkar et al. 2016), Web Questions (WebQ) (Berant et al. 2013), 2WikiMultiHopQA (2Wiki) (Ho et al. 2020), CoQA (Reddy, Chen, and Manning 2019), MS MARCO (Nguyen et al. 2016), PopQA (Mallen et al. 2023), Fermi (Kalyan et al. 2021), Musique (Trivedi et al. 2022), and Bamboogle (Press et al. 2023). These datasets are publicly available at HuggingFace or their official websites. To evaluate the generalizability of the methods, we select PopQA, Fermi, Musique, and Bamboogle as held-out datasets. We mix the training set of all remaining datasets for training. For all datasets, we prioritize the use of test sets for evaluation purposes. In cases where the test set is not available, we utilize the development set instead. It is worth noting that, though some datasets have provided golden reference passages for the answer, we do not use them in our experiment but use the passages retrieved from the retrieval sets in both training and inference stages, which aligns with practical applications. Exact match (EM) and F1 score are employed as evaluation metrics.
@@
For the retrieval sets, we follow previous studies (Yoran et al. 2023) and use the combination of Wikipedia and MS MARCO datasets as the retrieval corpus. Wikipedia contains high-quality human knowledge, which is helpful for many knowledge-intensive tasks. MS MARCO contains a large amount of web pages, which can provide information necessary for curating some natural language questions. We use the datasets that have already been preprossed into passages and released on HuggingFace. [^1] Wikipedia passages: https://huggingface.co/datasets/Tevatron/wikipedia-nq-corpus. MS MARCO passages: https://huggingface.co/datasets/Tevatron/msmarco-passage-corpus. The Wikipedia set has 21M passages, while the MS MARCO set has 8M passages. More details are provided in Appendix.
@@
We use E5-large (Wang et al. 2022) as the main retriever in our experiments. The impact of other retrievers, i.e., BM25 (Robertson and Zaragoza 2009), BGE-base (Xiao et al. 2023), and E5-base, is studied in our further analysis. Among these retrievers, BM25 is a non-neural sparse retrieval algorithm, while others are neural-based dense retrievers. In general, dense retrievers perform better on several benchmarks (Muennighoff et al. 2023).
Baseline Methods
@@
We consider both the base and instruction fine-tuned versions of Mistral-7b, LLaMA-2-7b, and LLaMA-2-13b as the backbone models, and compare our SPRING with the following baselines.